OpenStudy (anonymous):

For the polynomial function ƒ(x)= −2x2 − 4x + 16, find all local and global extrema. A) The local and global extrema are: (0, -4), (0, 2) and (-1, 18). B) The only extrema point is (−1, 18). C) No global extrema exist. D) No local extrema exist.

OpenStudy (anonymous):

are you using calculus or algebra?

OpenStudy (misty1212):

lol "extrema" fancy word for the second coordinate of the vertex!

OpenStudy (anonymous):

Algebra

OpenStudy (misty1212):

to find the vertex, first coordinate is \(-\frac{b}{2a}\)

OpenStudy (misty1212):

in your case it is \[-\frac{-4}{2\times (-2)}=1\]

OpenStudy (misty1212):

oops i meant \(-1\)

OpenStudy (anonymous):

My answer is B

OpenStudy (anonymous):

Then, yeah, just pick out the vertex a la @misty1212 . Since the highest exponent is 2, it has to be a parabola and only has 1 extrema

OpenStudy (anonymous):

@misty1212 am I right?

OpenStudy (misty1212):

yes

OpenStudy (anonymous):

thank you

OpenStudy (misty1212):

yw

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]

OpenStudy (anonymous):

can yall help me